OpenStudy (anonymous):

A radio station had 258 tickets to a concert. They gave away 5 times as many tickets to listeners as to employees. How many tickets did they give away to employees? ---please answer--- ------------------

OpenStudy (anonymous):

Since they give away 2 values, x, and y, and we know that y is 5 times what x is, we know the whole number can be divided by 6 to get the number that they give away to employees. 258 = x + y 258 =x + 5x 258/6 =43

OpenStudy (anonymous):

Thank you! @darthspectrum